2. Evaluate which of the following options would be your best investment based solely on the yield to maturity criterion. Option #1: Purchase a $60,000 coupon bond with a 7.5% coupon rate selling for $58,800 and maturing in 12 years Option #2: Purchase a $60,000 discount bond selling for $34,100 and maturing in 12 years Option #3: Loan a reliable friend $60,000 with promised repayments of $11,925.19 at the end of year 4, $28,442.01 at the end of year 8, and $50,876.44 at the end of year 12. Assume the payments represent 1/6, 2/6, and 3/6 repayment of the original loan amount respectively.
Answer #1

Answer:

Option 1 is financially better based on YTM. See calculations below.

Workings:
Yield of Maturity of Option 1:

Years Cash Flow Discount Factor at 5% Present Value DF at 8% Present Value
0 -58800 1.0000         -58,800.00 1.0000         -58,800.00
1 to 8 4500                                6.4632           29,084.46      5.7466           25,859.88
8 60000 0.67684           40,610.36 0.54027           32,416.13
          10,894.82               -523.99

IRR = YTM = 5% + (10894.82)/(10894.82+523.99) x 3 = 7.86% [A more precise calculation can result in 7.85%, that is if we take 7% and 8% to interpolate.]

Yield to Maturity of Option 2:

Years Cash Flow Discount Factor at 5% Present Value DF at 4% Present Value
0 -34100 1.0000         -34,100.00 1.0000         -34,100.00
12 60000                                0.5568           33,410.25      0.6246           37,475.82
              -689.75             3,375.82

IRR = YTM = 4% + (3375.82)/(3375.82+689.75) x 1 = 4.83%

Yield to Maturity of Option 3:

Years Cash Flow Discount Factor at 5% Present Value DF at 4% Present Value
0 -60,000.00 1.0000         -60,000.00 1.0000         -60,000.00
4    11,925.19                                0.8227             9,810.88      0.8548           10,193.70
8    28,442.01                                0.6768           19,250.67      0.7307           20,782.30
12    50,876.44                                0.5568           28,329.91      0.6246           31,777.27
          -2,608.54             2,753.27

IRR = YTM = 4% + (2753.27)/(2753.27+2608.54) x 1 = 4.51%

Hope this helps.
Please leave a thumbs up.
Thank you!